# Minimal Models for RNA Simulations

### v1

@@ 4 TIS-CAT
A technical challenge in explicitly treating both divalent and monovalent ions in simulations is that it dramatically decreases the computational efficiency. To overcome this problem the TIS-CAT hybrid model was developed in which monovalent ions are implicitly treated using the DH potential but divalent cations are explicitly included (Figure 1b) 23. The TIS-CAT model is valid when the concentration of monovalent ions (typically $>$50 mM) is higher than the divalent ion concentration, which is normally in the millimolar range in the folding experiments. We developed a comprehensive theoretical model combining the Reference Interaction Site Model (RISM) 64; 65; 66, grounded in liquid state theory, with coarse-grained molecular simulations. The RISM theory enables detailed calculations of the potential of mean force (PMF), which is distance and the ionic concentration-dependent effective potential between divalent ions and the phosphate group (Figure 4a). The isotropic PMF has two distinct minima: one for inner-sphere coordination (direct binding to the phosphate) and another for outer-sphere coordination (binding through water mediation). This dual-mode binding is crucial for accurate prediction of the thermodynamics of ion-induced RNA folding 23.
@@
Simulations using the hybrid model demonstrated that divalent ions, particularly Mg2+, have a profound effect on RNA folding through both inner- and outer-sphere coordination with phosphate groups. In smaller RNAs, such as the Beet Western Yellow Virus (BWYV) pseudoknot, Mg2+ predominantly binds via outer-sphere interactions, where the ion remains hydrated and interacts indirectly with RNA through water molecules. In contrast, in larger RNAs like the 58-nucleotide ribosomal RNA fragment (58-nt rRNA, Figure 4b), there is a significant change toward inner-sphere binding, where Mg2+ directly interacts with phosphate groups after partially losing its hydration shell. This change is due to the increased electrostatic potential in the more densely packed phosphate regions of larger RNA molecules, which necessitates stronger binding to stabilize the folded structure.
@@
The simulations also highlighted the role of bridging interactions on RNA stability 67. These occur when a single divalent ion binds simultaneously to two or more phosphate groups, acting as a bridge that stabilizes the RNA’s tertiary structure. Bridging interactions are particularly important in larger RNAs where multiple phosphate groups are in proximity. The simulations showed that reducing the concentration of divalent ions reduces the likelihood of the bridging interactions, leading to RNA destabilization. For example, in the 58-nt rRNA, specific nucleotides were identified as key binding sites for Mg2+ , with significant accumulation of ions around them 23. These nucleotides are located at structurally important positions, and the Mg2+ ions in these regions are crucial for maintaining the integrity of the folded RNA. This finding is particularly important for understanding the stability of larger RNA molecules, such as ribozymes and ribosomal RNAs, where multiple regions of the RNA come into proximity during folding.
@@
The charge density of divalent cations plays a crucial role in determining their coordination mode with RNA. The high charge density of Mg2+ favors outer-sphere binding in most small and intermediate-sized RNAs because it maintains a stable hydration shell. However, in regions of highly negative electrostatic potential, such as in larger RNA molecules or deep pockets of compact structures, Mg2+ can partially dehydrate and engage in inner-sphere coordination, as in the fluoride riboswitch. In contrast, Ca2+ , which has a lower charge density and thus weaker interaction with its hydration layer, dehydrates more readily and binds directly to phosphate groups in the inner-sphere coordination, even in smaller RNAs. This difference in binding behavior is important for understanding how different divalent ions influence RNA folding and stability. The findings suggest that Mg2+ binds RNA in multiple ways depending on the RNA size and structure, while Ca2+ is more rigid in its preference for direct phosphate binding.
@@
Figure 4: (a) The effective potential between phosphate and Mg2+ used in the TIS-CAT model (red) is a combination of the PMF derived from the RISM theory for the short range (grey dashed line) interaction and the Debye-Hückel theory for the long range (cyan) potential. There are two minima corresponding to the direct (dehydrated) and water-mediated binding of Mg2+ to the phosphate. (b) Preferential interaction coefficient, $\Gamma$, depending on the bulk Mg2+ concentrations as a function of monovalent salt concentration. $\Gamma$ values calculated from TIS-CAT simulations (symbols)23 are in excellent agreement with the experimental data (dashed curves)68. (c) Three distinct conformations of guanidine-II riboswitch from TIS-CAT simulations at different Mg2+ concentrations (upper) and secondary structures indicating the positions of experimental FRET dies (lower) 18. (d) Comparison of smFRET efficiency distribution between the simulations and experiments.18 The images in (c and d) were taken from 18 under the CC-BY license © 2022 Fuks, Falkner, Schwierz and Hengesbach.
@@
Recently, single molecule FRET experiments and simulations using the TIS-CAT model were combined to detect three distinct states that are populated at different Mg2+ concentrations in guanidine-II riboswitch that controls translation (Figure 4c) 18. When guanidine is bound (unbound) translation is turned on (suppressed) 69. Remarkably, the calculated distributions of FRET efficiency as a function Mg2+ using the TIS-CAT model without any adjustable parameter were in excellent agreement with experiments (Figure 4d)18. The TIS-CAT simulations were used to determine the structures of the states that are populated at three values of Mg2+ concentration. Note that unlike in guanidine-II riboswitch it is only upon binding to the metabolite is translation is suppressed in the SAM riboswitch. Nevertheless, the CG models accurately predict the outcomes of experiments, albeit using different resolution for the RNA.
@@ 5 Concluding Remarks
Although coarse-grained models for nucleic acids were first introduced only twenty-five years ago 10, a great deal of progress has been made in the intervening years, spurred largely by advances in the experimental fronts (such as single molecule studies, cryo-EM, time-resolved SAXS experiments, and ion-counting experiments). While these developments, some of which are sketched here, are encouraging, frontiers in RNA biology are rapidly moving to problems that invariably require coming to grips with more interactions involving multiple proteins, DNA, and RNA. The prime example is ribosome assembly, which has been extensively studied experimentally 70; 71; 72, thus producing accurate data. Although there are combination of a variety of simulation methods and experiments have provided invaluable insights 73, it is fair to say that computational methods alone are currently not capable of tackling the dynamics of ribosome assembly or similarly complicated problems. We believe that innovations in this context are sorely needed.
Acknowledgments: This work was supported by a grant from the National Science Foundation (CHE 2320256) and the Welch Foundation through the Collie-Welch Chair (F-0019). NH was supported by JST PRESTO (JPMJPR22EA).

### v2

@@ 4 TIS-CAT
A technical challenge in explicitly treating both divalent and monovalent ions in simulations is that it dramatically increases the computational cost. For instance, simulating a 35-nucleotide-long pseudoknot RNA requires a system comprising 3,423 beads, 3,303 of which are for K+ and Cl- ions. Consequently, the majority of the simulation time is spent computing interactions between monovalent ions. To overcome this problem, the TIS-CAT hybrid model was developed in which monovalent ions are implicitly treated using the DH potential but divalent cations are explicitly included (Figure 1b) 31. The TIS-CAT model is valid when the concentration of monovalent ions (typically $>$50 mM) is higher than the divalent ion concentration, which is normally in the millimolar range in the folding experiments. In the TIS-CAT model, we used the Reference Interaction Site Model (RISM) 68, grounded in liquid state theory in conjunction with coarse-grained molecular simulations. The RISM theory enables the calculations of the potential of mean force (PMF), which is the effective potential between divalent ions and the phosphate group distance that depends on the ion concentration (Figure 4a). The isotropic PMF has two distinct minima: one for inner-sphere coordination (direct binding to the phosphate) and another for outer-sphere coordination (binding through water mediation). The dual-mode binding is crucial for accurate prediction of the thermodynamics of ion-induced RNA folding 31.
@@
Simulations using the hybrid model demonstrated that divalent ions, particularly Mg2+, have a profound effect on RNA folding through both inner- and outer-sphere coordination with phosphate groups. In smaller RNAs, such as the Beet Western Yellow Virus (BWYV) pseudoknot, Mg2+ predominantly binds via outer-sphere interactions, where the ion remains hydrated and interacts indirectly with RNA through water molecules. In contrast, in the 58-nucleotide ribosomal RNA fragment (58-nt rRNA, Figure 4b), there is a significant shift toward inner-sphere binding, where Mg2+ directly interacts with phosphate groups after partially losing its hydration shell69. This change is due to the increased electrostatic potential in the more densely packed phosphate regions of larger RNA molecules, which necessitates stronger binding to stabilize the folded structure.
@@
The simulations also highlighted the role of bridging interactions on RNA stability 69. These occur when a single divalent ion binds simultaneously to two or more phosphate groups, acting as a bridge that stabilizes the RNA’s tertiary structure. Bridging interactions are particularly important in larger RNAs where multiple phosphate groups are in proximity. The simulations showed that lowering the concentration of divalent ions reduces the likelihood of bridging interactions, leading to RNA destabilization. Individual nucleotides were identified as either forming or not forming interactions with Mg2+, even in early folding or intermediate structures, depending on the bulk Mg2+ concentration. For example, in the 58-nt rRNA, specific nucleotides were identified as key binding sites for Mg2+, with significant accumulation of ions around them 31. These nucleotides are located at structurally important positions, and the Mg2+ ions in these regions are crucial for maintaining the integrity of the folded RNA. This finding is particularly important for understanding the stability of larger RNA molecules, such as ribozymes and ribosomal RNAs, where multiple regions of the RNA come into proximity during folding.
@@
The charge density of divalent cations plays a crucial role in determining their coordination mode with RNA. The high charge density of Mg2+ favors outer-sphere binding in most small and intermediate-sized RNAs because it maintains a stable hydration shell 69. However, in regions of highly negative electrostatic potential, such as in larger RNA molecules or deep pockets of compact structures, Mg2+ can partially dehydrate and engage in inner-sphere coordination, as in the fluoride riboswitch. In contrast, Ca2+ , which has a lower charge density and thus weaker interaction with its hydration layer, dehydrates more readily and binds directly to phosphate groups in the inner-sphere coordination, even in smaller RNAs. This difference in binding behavior is important for understanding how different divalent ions influence RNA folding and stability. The findings suggest that Mg2+ binds RNA in multiple ways depending on the RNA size and structure, while Ca2+ is more rigid in its preference for direct phosphate binding.
@@
Figure 4: (a) The effective potential between phosphate and Mg2+ used in the TIS-CAT model (red) is a combination of the PMF derived from the RISM theory for the short range (grey dashed line) interaction and the Debye-Hückel theory for the long range (cyan) potential. There are two minima corresponding to the direct (dehydrated) and water-mediated binding of Mg2+ to the phosphate. (b) Preferential interaction coefficient, $\Gamma$, depending on the bulk Mg2+ concentrations as a function of monovalent salt concentration. $\Gamma$ values calculated from TIS-CAT simulations (symbols)31 are in excellent agreement with the experimental data (dashed curves)70. (c) Three distinct conformations of guanidine-II riboswitch from TIS-CAT simulations at different Mg2+ concentrations (upper) and secondary structures indicating the positions of experimental FRET dies (lower) 24. (d) Comparison of smFRET efficiency distribution between the simulations and experiments.24 The images in (c and d) were taken from 24 under the CC-BY license © 2022 Fuks, Falkner, Schwierz and Hengesbach.
@@
Recently, single molecule FRET experiments and simulations using the TIS-CAT model were combined to detect three distinct states that are populated at different Mg2+ concentrations in guanidine-II riboswitch that controls translation (Figure 4c) 24. When guanidine is bound (unbound) translation is turned on (suppressed) 71. Remarkably, the calculated distributions of FRET efficiency as a function of Mg2+ concentration using the TIS-CAT model without any adjustable parameter were in excellent agreement with experiments (Figure 4d)24. The TIS-CAT simulations were used to determine the structures of the states that are populated at three values of Mg2+ concentration. Note that unlike the guanidine II riboswitch, the SAM riboswitch does not suppress translation until it binds to a metabolite. Nevertheless, the CG models accurately predict the outcomes of experiments, albeit using different resolution for the RNA.
The oxRNA model, which is also based on TIS representation of RNA, has been used to calculate the melting profiles of hairpins and the MMTV pseudoknot 19; 72. The initial version 19 of oxRNA did not account for electrostatic interactions. As a result, ion dependence of RNA folding was not investigated. Nevertheless, the predictions of the two melting temperatures in the MMTV pseudoknot at 1M salt were reasonably accurate. Subsequently 72, electrostatic interactions were incorporated using the DH potential. Application to supercoiling of double stranded RNA produced good agreement with experiments.
@@ 5 Concluding Remarks
Since the introduction of coarse-grained models to study RNA folding mechanisms twenty years ago 11, a great deal of progress has been made, spurred largely by advances in the experimental frontiers (such as single molecule studies, cryo-EM, time-resolved SAXS and ion-counting experiments). Although these developments, some of which are sketched here, are encouraging, the frontiers of RNA biology are rapidly moving to problems that invariably require coming to grips with complex systems involving multiple proteins, DNA, and RNA. To address such challenges, there is a clear need for more sophisticated coarse-grained models capable of accurately describing the formation of multi-component complexes, such as RNA-protein assemblies, while remaining computationally applicable to large-scale systems. One example is ribosome assembly, which has been extensively studied experimentally 73; 74; 75. A normal mode analysis using elastic network representation provided insights into the dynamics of ribosomes in the ground state 76, and a single-bead coarse-grained simulation captured certain motions beyond the ground state 77. Nevertheless, it is fair to say that computational methods alone are currently not capable of tackling the dynamics of ribosome assembly or spliceosomes.
Acknowledgments: This work was supported by a grant from the National Science Foundation (CHE 2320256) and the Welch Foundation through the Collie-Welch Chair (F-0019). NH was supported by JST PRESTO (JPMJPR22EA). We thank Balaka Mondal for proofreading the manuscript and the two anonymous referees who provided valuable feedback to help us improve it.